1.    A solution is prepared by mixing 16.0 g of CH3OH and 50.0 g of water.

a.     What is the mass percent of methanol in the solution?

b.    What is the mole fraction of each component in the solution?

c.     If the density of solution is 0.955 g/mL, what is the molarity and molality of CH3OH, respectively, in the solution?


1
Expert's answer
2021-10-28T02:59:04-0400

a) 16/(16+50)*100=24.24%

b) v(ch3oh)=16/32=0.5mol

v(h2o)=50/18=2.78mol

c) 0.955*32=30.56